OpenStudy (anonymous):

Which facts are true for the graph of the function below? Check all that apply. f(x)=(2/5)^x A. The range of F(x) is y > 0. B. The domain of F(x) is x > 0. C. The y-intercept is (0, 1). D. It is increasing. E. It is decreasing. F. The x-intercept is (1, 0).

OpenStudy (anonymous):

this graph is also decreasing..

jagr2713 (jagr2713):

A is false. ¯¯¯¯¯¯¯¯¯ The x-intercept occurs where y = 0, but y does not reach 0,

jagr2713 (jagr2713):

B is false. The y-intercept occurs where x = 0, so y = (2/5)^(0) Any number raised to the 0th power is one, so y-int. (0, 1)

jagr2713 (jagr2713):

C is true. If the largest exponent in the numerator is equal to the largest exponent in the denominator, the horizontal asymptote is the coefficient of the variable with the largest exponent in the numerator divided by the coefficient of the variable with the largest exponent in the denominator 2 / 5 = 0.1 Range: y > 0.1

jagr2713 (jagr2713):

E is true and D is false. f(x) is a fraction and, as such, the denominator cannot equal zero, lest division by zero occurs, which is undefined, 5^x ≠ 0 Any number raised to a power cannot equal zero, so there are no constraints, so Domain: All Real x,

jagr2713 (jagr2713):

False, when x equals 1, y equals 2/5

OpenStudy (anonymous):

what about F?

jagr2713 (jagr2713):

F is False, when x equals 1, y equals 2/5

OpenStudy (anonymous):

A is true.

OpenStudy (anonymous):

The base of an exponential function cannot be a negative number. A. True B. False

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

true

OpenStudy (anonymous):

Which best describes the asymptote of an exponential function of the form F(x) = b^x? A. Horizontal asymptote at y = 1 B. Horizontal asymptote at y = 0 C. Vertical asymptote at x = 0 D. Vertical asymptote at x = 1

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

last one then i gtg for a 20mins

OpenStudy (anonymous):

ok thanks!

jagr2713 (jagr2713):

i gtg sorry brb later